OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

yoooooo

Nnesha (nnesha):

y = mx +b slope intercept form where m is slope so replace m by -6

OpenStudy (ribhu):

y-y1 = m(x-x1)

OpenStudy (anonymous):

hi @iambatman

OpenStudy (anonymous):

ok @Nnesha

OpenStudy (anonymous):

this is point slope form, so you coordinates are (x1,y1) just use the equation ribhu showed you, and slope is m.

OpenStudy (ribhu):

now x1,y1 is the point from which the line passes and slope of line is m.

OpenStudy (ribhu):

@heymydearpeople try to do on your own using this concept.

OpenStudy (anonymous):

@Nnesha

OpenStudy (ribhu):

its not slope intercept form @Nnesha

Nnesha (nnesha):

yes ??

OpenStudy (anonymous):

which form is correct?

Nnesha (nnesha):

your answer are in point slope form so use that one

OpenStudy (anonymous):

the one that @iambatman said

OpenStudy (anonymous):

and @ribhu

Nnesha (nnesha):

yes

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1}) \]

OpenStudy (anonymous):

here is the part that always confuses me whatt do i plug in @Nnesha @iambatman

Nnesha (nnesha):

(x . y ) replace x1 and y1

OpenStudy (anonymous):

You have the points (-2, -3) this is your (x1,y1) and your slope is the m in the equation which is -6.

OpenStudy (anonymous):

what is x and y

OpenStudy (anonymous):

You leave that alone

OpenStudy (anonymous):

? @iambatman

OpenStudy (anonymous):

You don't plug anything in the y and x, just y_1 and x_1 even look at your choices.

OpenStudy (anonymous):

is it a ? @iambatman

Nnesha (nnesha):

yes :)

OpenStudy (anonymous):

I agree as well :)